Evaluate 7i4-12i squared: Evaluate 7i4-12i squared
Answer by ankor@dixie-net.com(4532)

Evaluate 7i^4 - 12i^2
:
we know that i^2 = -1
and i^4 = i^2*i^2 therefore i^4 = -1 * -1 = +1
:
we have
7(1) - (12*-1)
7 - (-12)
7 + 12 = 19